Skip to content

RAMAN Setup

White edited this page Mar 21, 2016 · 2 revisions

RAMAN Setup

21 March 16:

  • Room booked at Base for spectro Setup
  • 500g of bee pollen sample received for bulk testing
  • Laser and spectro setup OK
  • Two lasers to be tested asap

Clone this wiki locally